• Main INDEX
  • Monthly INDEX
  • PREV

    User name fansler

    Log entry time 11:44:22 on February 20, 2008

    Entry number 228569

    keyword=Q1 liquid level readback

    Both Q1 He liquid level readbacks have been converted from 4-20ma to 0-10V, the noise on the left Q1 is gone, and all offsets between cryo and Hall A signals have been corrected.

    A copy of this log entry has been emailed to: fansler@jlab.org